40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 9 04:26:51 2012, Sasha, Update, Simulations, All c1spx screens working 
    Reply  Thu Aug 9 07:24:58 2012, Sasha, Update, Simulations, All c1spx screens working Screenshot.pngScreenshot-1.png
Message ID: 7133     Entry time: Thu Aug 9 07:24:58 2012     In reply to: 7132
Author: Sasha 
Type: Update 
Category: Simulations 
Subject: All c1spx screens working 

Quote:

As the subject states, all screens are working (including the noise screens), so we can keep track of everything in our model! :D I figured out that I was just getting nonsense (i.e. white noise) out of the sim plant cause the filter matrix (TM_RESP) that controlled the response of the optics to a force (i.e. outputted the position of the optic DOF given a force on that DOF and a force on the suspension point) was empty, so it was just passing on whatever values it got based on the coefficients of the matrix without DOING anything to them. In effect, all we had was a feedback loop without any mechanics.

I've been working on getting the mechanics of the suspensions into a filter/transfer function form; I added something resembling that into foton and turned the resulting filter on using the shiny new MEDM screens. However, the transfer functions are a tad wonky (particularly the one for pitch), so I shall continue working on them. It had a dramatic effect on the power spectrum (i.e. it looks a lot more like it should), but it still looks weird.

Still haven't found the e-log Jamie and Rana referred me to, concerning the injection of seismic noise into the simulation. I'm not terribly worried though, and will continue looking in the morning. Worst case scenario, I'll use the filters Masha made at the beginning of the summer.

Masha and I ate some of Jamie's popcorn. It was good.

 Okay! Attached are two power spectra. The first is a power spectrum of reality, the second is a power spectrum of the simPlant. Its looking much better (as in, no longer obviously white noise!), but there seems to be a gain problem somewhere (and it doesn't have seismic noise). I'll see if I can fix the first problem then move on to trying to find the seismic noise filters.

Attachment 1: Screenshot.png  196 kB  | Hide | Hide all
Screenshot.png
Attachment 2: Screenshot-1.png  201 kB  | Hide | Hide all
Screenshot-1.png
ELOG V3.1.3-